Circular No. 7521

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2000 S 5 AND S/2000 S 6 A team led by B. Gladman (cf. IAUC 7512) reports the discovery of

two more satellites of Saturn (selected CCD astrometry below). Separate computations by the undersigned and by W. J. Gray (Bowdoinham, ME) indicate i = 49^{o}, e = 0.14, a = 0.074 AU and 47^{o}, 0.38, 0.085 AU, respectively, for the pair. S/2000 S 2 (cf. MPC 41527) has i = 48^{o}, e = 0.46, a = 0.098 AU, whereas S/2000 S 1 and Saturn IX (Phoebe) have retrograde orbits (i = 173^{o}). The orbits of S/2000 S 3 and S 4 are very indeterminate, and further observations are urgently needed (cf. IAUC 7513).

2000 UT α_{2000} δ_{2000} R Observer
Aug. 7.39963 3^{h}52^{m}08.^{s}02 +18^{o}20′03′′.0 22.6 Gladman
Sept. 24.44657 3 57 48.21 +18 27 23.0  Kavelaars
Nov. 2.27282 3 50 26.67 +17 50 21.9  Hergenrother
3.29969 3 50 08.33 +17 49 01.7  _{”}
4.22350 3 49 51.61 +17 47 50.0  UT1 Team
Sept. 23.44526 3 57 45.16 +17 42 52.4 22.8 Kavelaars
24.46410 3 57 41.36 +17 42 34.8 22.6  _{”}
Nov. 2.29914 3 50 17.20 +17 21 08.1  Hergenrother
3.30891 3 49 59.47 +17 20 24.2  _{”}
4.23929 3 49 42.93 +17 19 44.5 22.9 UT1 Team

B. Gladman (European Southern Obs.). 2.2-m refl. Measurer J.-M. Petit. J. Kavelaars and B. Gladman (Mauna Kea). 3.6-m Canada-France-Hawaii

Telescope. Measurer Gladman. C. W. Hergenrother, R. J. Whiteley and D. Means (Steward Observatory,

Kitt Peak). 2.3-m reflector. Measurer Hergenrother. UT1 Science Team (Paranal). ESO VLT-UT1 8-m refl. Measurer Gladman.

2000TT α_{2000} δ_{2000} Δ r_{S} ϵ Δα′′ Δδ′′
Nov. 12 3^{h}47.^{m}42 +17^{o}37.^{′}4 8.120 0.072 171.^{o}0 1638+ 546+
22 3 44.11 +17 23.6 8.104 0.073 176.8 1659+ 341+
Dec. 2 3 40.80 +17 10.1 8.121 0.075 166.6 1634+ 127+
12 3 37.68 +16 57.6 8.169 0.076 155.8 1569+ 91−
22 3 34.91 +16 46.9 8.246 0.078 145.0 1469+ 304−
Nov. 12 3 47.33 +17 14.0 8.180 0.080 171.0 1562+ 855−
22 3 44.09 +17 06.9 8.164 0.076 176.5 1642+ 663−
Dec. 2 3 40.85 +17 00.4 8.179 0.072 166.6 1678+ 456−
12 3 37.78 +16 55.1 8.224 0.068 155.8 1660+ 240−
22 3 35.05 +16 51.6 8.297 0.064 145.0 1585+ 23−
2000 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden